# Bourbaki degree of pairs of projective surfaces

### v2

@@ Acknowledgments
FM is financed by the São Paulo Research Foundation (FAPESP), Brasil. Process Number #2021/10550-4. FM is currently a PhD. student under the advice of Marcos Jardim and Daniele Faenzi, with partial funding by the Bridges ”Brazil-France interplays in Gauge Theory, extremal structures and stability” projects ANR-21-CE40-0017 and ANR-17-EURE-0002. We thank Marcos Jardim, Daniele Faenzi and Alan Muniz for fruitful discussions and suggestions regarding this work.
2 Basic setting and $m$-syzygy sequences on $\mathbb{P}^{n}$
We work over an algebraically closed field $\kappa$ of characteristic zero, although most of the definitions can be considered under relaxed assumptions. We denote by $R=\kappa[x_{0},\ldots,x_{n}]$ the graded ring of polynomials over $\kappa$ and $\mathbb{P}^{n}\doteq\mathbb{P}^{n}_{\kappa}=\operatorname{Proj}R$ the projective $n$ space over $\kappa$.
In this section, we recall from [11] the definition of the logarithmic tangent sheaf associated to a sequence of $k$ polynomials, and introduce some notation inspired by the literature about the case $k=1$.
For a sequence $\sigma=(f_{1},\ldots,f_{k})$ of homogeneous polynomials with degrees $d_{1}+1,\ldots,d_{k}+1$, we consider the Jacobian matrix $\nabla\sigma$ as a morphism between locally free sheaves:
$\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}\xrightarrow{\nabla\sigma}\bigoplus_{i=1}^{k}\mathcal{O}_{\mathbb{P}^{n}}(d_{i})$
so the kernel sheaf ${\mathcal{T}}_{\sigma}\doteq\ker(\nabla\sigma)$, is called the logarithmic tangent sheaf associated to $\sigma$. We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$ the cokernel sheaf and by $\mathcal{M}_{\sigma}\doteq\operatorname{im}(\nabla\sigma)$ the image sheaf. We denote by $d\doteq\sum_{i=1}^{n}d_{i}$ the total degree of a sequence $\sigma$. Assuming the generic rank of $\operatorname{im}(\nabla\sigma)$ is $k$, it follows that ${\mathcal{T}}_{\sigma}$ is a reflexive sheaf of rank n+1kn+1-k.
2.1 Normal sequences
By construction, the reduced support of the sheaf $\mathcal{Q}\sigma$ coincides with the vanishing locus $\Xi\sigma\doteq V(\bigwedge^{k}\nabla\sigma)$ of the $k\times k$-minors of the matrix $\nabla\sigma$. Assuming $\sigma$ is algebraically independent, the generic rank of $\nabla\sigma$ is $k$, so that $\operatorname{{codim}}\Xi_{\sigma}\geq 1$.
Moreover, since $\mathcal{Q}\sigma$ is a torsion sheaf, the condition $c_{1}(\mathcal{Q}\sigma)=0$ is equivalent to $\operatorname{{codim}}(\operatorname{supp}(\mathcal{Q}_{\sigma}))\geq 2$. In terms of the matrix $\nabla\sigma$, $\operatorname{{codim}}\Xi_{\sigma}\geq 2$ if and only if the greatest common divisor among all $(k\times k)$-minors is one. When $k=1$, we have
$\operatorname{{codim}}(Z_{f})=\operatorname{{codim}}(\mathcal{Q}_{f})\geq 2\iff V(f)\text{ is normal}.$
Inspired by the behavior above, we call a sequence $\sigma$ normal when $\operatorname{{codim}}(\operatorname{supp}(\mathcal{Q}_{\sigma}))\geq 2$.
We note that the two schemes $\operatorname{supp}(\mathcal{Q}_{\sigma})$ and $\Xi_{\sigma}$ may have different scheme structures and, in particular, different degrees. On one hand, the scheme structure of the support $\operatorname{supp}(\mathcal{Q}_{\sigma})=V(\operatorname{Ann}\mathcal{Q}_{\sigma})$ is induced by the annihilator ideal sheaf of $\mathcal{Q}_{\sigma}$, defined locally as the annihilator ideal of the associated module. On the other hand, the scheme structure of $\Xi_{\sigma}$ is induced locally by the 0-th Fitting ideal of the $\mathcal{Q}_{\sigma}$. These schemes have the same reduced locus, since the support of these two ideals coincide, but in general $\operatorname{Fitt}_{0}\subseteq\operatorname{Ann}$ (see, for example, [9, Proposition 20.7]).
For $k=1$, when $\sigma=f$, we have an isomorphism $\mathcal{Q}f\simeq\mathcal{O}{Z_{f}}$. This holds because $\mathcal{M}f\simeq\mathcal{I}{Z_{f}}(d)$ is the ideal sheaf of the scheme $Z_{f}$, and whenever $f$ is reduced, the schemes $\Xi_{f}$ and $\operatorname{supp}(\mathcal{Q}{\sigma})$ coincide.
If $\sigma$ is a normal sequence, then the Hilbert polynomial of the cokernel sheaf $\mathcal{Q}_{\sigma}$ is of the form
$\operatorname{Hilb}(\mathcal{Q}_{\sigma},t)=\frac{m(\sigma)}{(n-2)!}t^{n-2}+\ldots$
where $m(\sigma)\geq 0$ is a discrete invariant so that $m(\sigma)=0$ if and only if $\operatorname{{codim}}(\mathcal{Q}_{\sigma})\geq 3$. However, $m(\sigma)$ generally coincides with neither the schematic degree of $\Xi_{\sigma}$ nor that of $\operatorname{supp}(\mathcal{Q}_{\sigma})$, as we explore in Example 2 at the end of this section.
2.2 $m$-syzygy sequences and exponents
We may consider a minimal free resolution of the sheaf ${\mathcal{T}}_{\sigma}$, which will be given in the following form:
$\ldots\rightarrow\bigoplus_{i=1}^{m}\mathcal{O}_{\mathbb{P}^{n}}(-e_{i})\xrightarrow{\rho}{\mathcal{T}}_{\sigma}\rightarrow 0.$
In which case, we say $\sigma$ is an $m$-syzygy sequence.
We note that $m\geq n$ and $\sigma$ is said to be free if and only if ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles, or, equivalently, when $m=n$.
The integers $e_{1}\leq\ldots\leq e_{m}$ will be called exponents, and the first $\operatorname{indeg}({\mathcal{T}}_{\sigma})\doteq e_{1}=e$ will be called the initial degree of the sheaf ${\mathcal{T}}_{\sigma}$. This is the minimum integer $e\geq 0$ such that there is a non-trivial syzygy $\nu$ of degree $d$ for the Jacobian matrix $\nabla\sigma$. Since there is an injection ${\mathcal{T}}_{\sigma}\hookrightarrow\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}$ and the latter is a $\mu$-semistable sheaf, all exponents satisfy $e_{i}\geq 0$.
A sequence $\sigma=(f_{1},\ldots,f_{k})$ is called compressible if, after a linear change of variables, there is a variable that does not occur in any of the $f_{1},\ldots,f_{k}$. The number of variables omitted in this way will give rise to syzygies of degree zero for $\nabla\sigma$, so the following hold (for proofs, see [11, Lemma 2.7 and Lemma 2.8]):
@@ 2.2 $m$-syzygy sequences and exponents
The following quantities coincide:
  • The number of variables omitted in $f_{1},\ldots,f_{k}$;
  • The discrete invariant $h^{0}({\mathcal{T}}_{\sigma})$;
  • The number of 0’s in a vector of exponents for ${\mathcal{T}}_{\sigma}$.
In particular, $e=0$ if and only if $\sigma$ is compressible.
For $n=3$ and when $\sigma=(f,g)$ is a normal sequence, we show the conditions above are equivalent to $m(\sigma)$ attaining its maximum value, see Section 3.1.
@@
so both schemes $\Xi_{\sigma}$ and $\operatorname{supp}(\mathcal{Q}_{\sigma})$ are non-reduced, with degrees $4$ and $6$, respectively. However, from the formula for $\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})$ we conclude $m(\sigma)=5$.
3 The Bourbaki degree of pairs of projective surfaces
In this section, we develop the concept of the Bourbaki degree of pairs of projective surfaces on $\mathbb{P}^{3}$, determined by normal sequences $\sigma=(f,g)$ of homogeneous polynomials. We start with definitions and first results (3.1), followed by a reduction to the case of a projective plane curve (3.2) and finish with comparison results relating the geometry of the Bourbaki scheme and the associated logarithmic sheaf, using free resolutions (3.3), in particular introducing the class of nearly-free sequences $\sigma$, which are characterized by $\operatorname{Bour}(\sigma)=1$.
3.1 Framework and first results
By a sequence $\sigma=(f,g)$, unless otherwise stated, we mean an algebraically independent sequence of two homogeneous polynomials in $R\doteq\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1$, $\deg(g)=d_{g}+1$. We also work only with normal sequences, as defined previously in 2.1. By curve we mean a locally Cohen-Macaulay closed subscheme of $\mathbb{P}^{3}$ of pure dimension one.
and we denote the kernel by ${\mathcal{T}}_{\sigma}$, the image by $\mathcal{M}_{\sigma}$ and the cokernel by $\mathcal{Q}_{\sigma}$. We say a sequence $\sigma$ is free whenever ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles, following [11].
Lemma 3.
Proof.
@@ 3.1 Framework and first results
where $B_{\nu}\subset\mathbb{P}^{3}$ is the curve associated by $\nu$, described by the Serre correspondence (see [14, Theorem $4.1$]). Moreover, we have $\deg(B_{\nu})=c_{2}({\mathcal{T}}(e))$. Since ${\mathcal{T}}_{\sigma}\hookrightarrow{\mathcal{O}}_{\mathbb{P}^{3}}^{\oplus 4}$ and the latter is a $\mu$-semistable sheaf, $e\geq 0$. The following proposition describes a formula for the degree $\deg(B_{\nu})$ in terms of the discrete invariants $e$, $d_{f}$, $d_{g}$ and $m(\sigma)$.
@@
We define the Bourbaki degree of a sequence $\sigma$, inspired by [17, Definition 2.4].
@@
It follows from construction that the sequence $\sigma$ is free if and only if $\operatorname{Bour}(\sigma)=0$.
@@
From [13, Proposition 4.1], we have the formula
@@
Theorem A. Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $\kappa[x_{0},\ldots,x_{3}]$, with degrees $d_{f}+1,d_{g}+1$ respectively. Then:
@@
For $(c)$, we start by pointing out that the equivalence $(2)\iff(3)$ is the content of [11, Lemma 2.7]. Moreover, $(3)\Rightarrow(4)$ using that $c_{1}({\mathcal{T}}_{\sigma})=-d$ and [11, Lemma 2.8]. The implication $(4)\Rightarrow(1)$ can be obtained using the Bourbaki degree formula, since $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$ and $\operatorname{Bour}(\sigma)=0$, as $\sigma$ is free.
@@
The bound obtained in 3.1, $(b)$ for $m(\sigma)$ relates to a known bound in foliation theory. We will use that ${\mathcal{T}}_{\sigma}(1)$ is the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$ of degree $d=d_{f}+d_{g}$ (see [11, Section 9]). Let $C$ be the one-dimensional part of the singular scheme of this foliation. From the formulas of discrete invariants in [2, Theorem 3.1],
@@
and therefore $\deg(C)\leq d^{2}+d+1$, a bound that can be found in more generality for foliations in [20, Corollary 4.8].
@@
    $\operatorname{Bour}(\sigma)\leq m_{0}$, and equality holds if and only if $m(\sigma)=0$ and $e=d$.
@@
The claim $(a)$ follows from the formula $\operatorname{Bour}(\sigma)\geq 0$, since $m(\sigma)\geq 0$ and $e(e-d)\leq 0$, from the inequality $e\leq d$. This also shows equality occurs whenever e(ed)e(e-d) and $m(\sigma)$ are both zero, therefore $e=d$, since $e=0$ means $\operatorname{Bour}(\sigma)=0$ from compressibility. The converse also follows simply from the formula.
@@
If $\sigma=(f,g)$ is a pencil of surfaces of the same degree, say $d_{f}=d_{g}=p$, then $m(\sigma)=0$ if $\sigma$ is a regular pencil and, moreover, every singular member $V(z_{0}f+z_{1}g)\subset\mathbb{P}^{3}$ is normal. This follows from the description
$\operatorname{supp}(\mathcal{Q}_{\sigma})_{\text{red}}=\bigcup_{[z_{0}:z_{1}]\in\mathbb{P}^{1}}\operatorname{Sing}(V(z_{0}f+z_{1}g))$
obtained in [11, Lemma 2.17], since normal projective surfaces have isolated singularities and regular sequences have only a finite number of singular members, therefore $\dim(\operatorname{supp}(\mathcal{Q}_{\sigma}))=0$.
@@
Here, $m(\sigma)=0$, $\operatorname{Bour}(\sigma)=12$ and $c_{3}({\mathcal{T}}_{\sigma})=32$. Moreover, ${\mathcal{T}}_{\sigma}$ admits a locally free resolution of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
Remark 12.
We have not found any examples of sequences with $m(\sigma)=0$ so that $e<d$, which motivates the conjecture: does $m(\sigma)=0$ implies $e=d$?
3.2 The Bourbaki degree of a plane curve from the Bourbaki degree of a pair of surfaces
In this section, we obtain the formula of the Bourbaki degree of a projective plane curve $X=V(g)\subset\mathbb{P}^{2}$, where $g\in\kappa[x_{0},x_{1},x_{2}]$ is a reduced homogeneous polynomial, introduced in [17], as a special case of our formula, by considering the pair $\sigma=(x_{3},g)$ in $\mathbb{P}^{3}$.
Let $g\in\kappa[x_{0},x_{1},x_{2}]$ be a square-free polynomial, so that the projective curve $X=V(g)\subset\mathbb{P}^{2}$ is reduced with isolated singularities, and the algebraically independent sequence $\sigma=(x_{3},g)$, denoting the plane $H=V(x_{3})\simeq\mathbb{P}^{2}$. We consider $S=V(g)\subset\mathbb{P}^{3}$ as a projective surface whose singular locus consists of possibly non-reduced lines, with the same multiplicity at each corresponding intersection point with $H$.

### v3

@@ Acknowledgments
FM is financed by the São Paulo Research Foundation (FAPESP), Brasil. Process Number #2021/10550-4. FM is currently a PhD. student under the advice of Marcos Jardim and Daniele Faenzi, with partial funding by the Bridges ”Brazil-France interplays in Gauge Theory, extremal structures and stability” projects ANR-21-CE40-0017 and ANR-17-EURE-0002. We thank Marcos Jardim, Daniele Faenzi, Alan Muniz, Victor Cordeiro and Abbas Nejad for fruitful discussions and suggestions regarding this work.
2 The Bourbaki degree of pairs of projective surfaces
In this section, we develop the concept of the Bourbaki degree of pairs of projective surfaces on $\mathbb{P}^{3}$, determined by normal sequences $\sigma=(f,g)$ of homogeneous polynomials. We start with definitions and first results (Section 2.1), followed by a reduction to the case of a projective plane curve (Section 2.2) and finish with comparison results relating the geometry of the Bourbaki scheme and the associated logarithmic sheaf, using free resolutions (Section 2.3), in particular introducing the class of nearly free sequences $\sigma$, which are characterized by $\operatorname{Bour}(\sigma)=1$.
2.1 Framework and first results
By a sequence $\sigma=(f,g)$, unless otherwise stated, we mean an algebraically independent sequence of two homogeneous polynomials in $R\doteq\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1$, $\deg(g)=d_{g}+1$. By curve we mean a locally Cohen-Macaulay closed subscheme of $\mathbb{P}^{3}$ of pure dimension one.
and we denote the kernel by ${\mathcal{T}}_{\sigma}$, the image by $\mathcal{M}_{\sigma}$ and the cokernel by $\mathcal{Q}_{\sigma}$. We say a sequence $\sigma$ is free whenever ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles, following [undefj].
@@ 2.1 Framework and first results
Proof.
As in the introduction, the Jacobian scheme of $\nabla\sigma$ is defined as
$\Xi_{\sigma}\doteq V\left(\bigwedge^{2}\nabla\sigma\right)=V(\operatorname{Fitt}_{0}(\nabla\sigma))$
the zero-locus of the $2\times 2$ minors of $\nabla\sigma$ or, in other words, the $0$-th Fitting scheme of $\nabla\sigma$. From the general theory of Fitting ideals (see, for example, [undefh, Chapter 20]), the annihilator ideal of $\mathcal{Q}_{\sigma}$ contains the Fitting ideal, and they have the same support, namely
$\operatorname{supp}(\mathcal{Q}_{\sigma})_{\text{red}}=V(\operatorname{Fitt}_{0}(\nabla\sigma))$
as closed subsets, and therefore the irreducible components of the schematic support $\operatorname{supp}(\mathcal{Q}_{\sigma})$ coincide with the irreducible components of $\Xi_{\sigma}$. In general, however, their schematic structure may be different, as we explore in the following example. Denote by $(\Xi_{\sigma})_{1}$ the one-dimensional part of the scheme $\Xi_{\sigma}$.
@@
so both schemes $\Xi_{\sigma}$ and $\operatorname{supp}(\mathcal{Q}_{\sigma})$ are non-reduced, with degrees $4$ and $6$, respectively. From the formula for $\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})$, since ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$ we conclude $m(\sigma)=5$. The right approach to computing $m(\sigma)$ in general is using the associativity formula, as we will in the next examples.
@@
where $B_{\nu}\subset\mathbb{P}^{3}$ is the curve associated by $\nu$, described as the Serre correspondence (see [undefn, Theorem $4.1$]) between the sheaf ${\mathcal{T}}_{\sigma}$ and the scheme $B$. Moreover, we have $\deg(B_{\nu})=c_{2}({\mathcal{T}}(e))$. Since ${\mathcal{T}}_{\sigma}\hookrightarrow{\mathcal{O}}_{\mathbb{P}^{3}}^{\oplus 4}$ and the latter is a $\mu$-semistable sheaf, $e\geq 0$. The following proposition describes a formula for the degree $\deg(B_{\nu})$ in terms of the discrete invariants $e$, $d_{f}$, $d_{g}$ and $m(\sigma)$.
@@
We define the Bourbaki degree of a sequence $\sigma$, inspired by the construction in [undefr, Definition 2.4].
@@
It follows from construction that the sequence $\sigma$ is free if and only if $\operatorname{Bour}(\sigma)=0$, since every short exact sequence of the form
$0\to\mathcal{O}_{\mathbb{P}^{3}}(-e)\to{\mathcal{T}}_{\sigma}\to\mathcal{O}_{\mathbb{P}^{3}}(e-d)\to 0$
splits.
@@
From [undefm, Proposition 4.1], we have the formula
@@
Theorem A.
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $\kappa[x_{0},\ldots,x_{3}]$, with degrees $d_{f}+1,d_{g}+1$ respectively. Then:
@@
For $(c)$, we start by pointing out that the equivalence $(2)\iff(3)$ is the content of [undefj, Lemma 2.7]. Moreover, $(3)\Rightarrow(4)$ using that $c_{1}({\mathcal{T}}_{\sigma})=-d$ and [undefj, Lemma 2.8]. The implication $(4)\Rightarrow(1)$ can be obtained using the Bourbaki degree formula, since $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$ and $\operatorname{Bour}(\sigma)=0$, as $\sigma$ is free.
@@
The bound obtained in Theorem A, $(b)$ for $m(\sigma)$ relates to a known bound in foliation theory. We will use that ${\mathcal{T}}_{\sigma}(1)$ is the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$ of degree $d=d_{f}+d_{g}$ (see [undefj, Section 9]). Let $C$ be the one-dimensional part of the singular scheme of this foliation. From the formulas of discrete invariants in [undefa, Theorem 3.1] given for generic foliations,
@@
and therefore $\deg(C)\leq d^{2}+d+1$, a bound that can be found in more generality for foliations in [undefv, Corollary 4.8].
@@
    $\operatorname{Bour}(\sigma)\leq m_{0}$, and equality holds if any of the equivalent facts hold:
    • (1)
      $m(\sigma)=0$;
    • (2)
      A minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
      $0\to\mathcal{O}_{\mathbb{P}^{3}}(-d-d_{1})\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d-d_{2})\xrightarrow{\varphi}\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}(-d)\xrightarrow{\psi}{\mathcal{T}}_{\sigma}\to 0,$
      a Buchsbaum–Rim resolution, where
      $\psi=\begin{pmatrix}0&\Delta_{12}&\Delta_{13}&\Delta_{14}\\ -\Delta_{12}&0&\Delta_{23}&\Delta_{24}\\ -\Delta_{13}&-\Delta_{23}&0&\Delta_{34}\\ -\Delta_{14}&-\Delta_{24}&-\Delta_{34}&0\end{pmatrix},\quad\phi=\begin{pmatrix}\partial_{0}f&-\partial_{0}g\\[2.0pt] -\partial_{1}f&\partial_{1}g\\[2.0pt] \partial_{2}f&-\partial_{2}g\\[2.0pt] -\partial_{3}f&\partial_{3}g\end{pmatrix}$,
      and each $\Delta_{ij}$ is the $(i,j)$-minor of the matrix $\nabla\sigma$. In particular, $e=d$.
@@
The claim $(a)$ follows from the formula $\operatorname{Bour}(\sigma)\geq 0$, since $m(\sigma)\geq 0$ and $e(e-d)\leq 0$, from the inequality $e\leq d$. This also shows equality occurs whenever e(ed)e(e-d) and $m(\sigma)$ are both zero, therefore $e=d$, since $e=0$ means $\operatorname{Bour}(\sigma)=0$ from compressibility. The converse also follows simply from the formula, assuming that $m(\sigma)=0$ and $e=d$.
To show the equivalences above, we start by noting that $m(\sigma)=0$ implies that the grade of the ideal of $(2\times 2)$-minors is three, and therefore the Buchsbaum–Rim complex is a minimal free resolution for $\nabla\sigma$. From this fact, we obtain that $e=d$. For the converse, since $e=d$ and the Buchsbaum–Rim complex is a minimal resolution, we obtain that the grade of the ideal of minors of $\nabla\sigma$ is at least three, hence $\dim(\mathcal{Q}_{\sigma})=0$ and thus $m(\sigma)=0$.
@@
Let $\sigma=(f,g)$ be a sequence of surfaces of the same degree, say $d_{f}=d_{g}=p$, called a pencil of surfaces. It defines a 11-parameter family of projective surfaces as:
$V_{\sigma}\doteq\{V(z_{0}f+z_{1}g)\subset\mathbb{P}^{3}:z=[z_{0}:z_{1}]\in\mathbb{P}^{1}\},$
and each hypersurface $V(z_{0}f+z_{1}g)$ is called a member of the pencil $V_{\sigma}$. The set where each surface is singular is the vanishing locus of a polynomial $p=p(z)$ (the discriminant) inside $\mathbb{P}^{1}$, so it must be either a finite subset or the whole family. When this set is finite, the pencil is called a regular pencil, otherwise is called irregular.
In [undefj, Lemma 2.17], the authors describe the support of the sheaf $\mathcal{Q}_{\sigma}$ for pencils of surfaces as the union of singular loci:
$(\Xi_{\sigma})_{\text{red}}=\bigcup_{[z_{0}:z_{1}]\in\mathbb{P}^{1}}\operatorname{Sing}(V(z_{0}f+z_{1}g)).$
Hence, whenever a normal sequence $\sigma=(f,g)$ is a regular pencil where every singular member has isolated singularities, we conclude $\dim(\Xi_{\sigma})=0$ from the description above, and thus $m(\sigma)=0$.
@@
Here, $m(\sigma)=0$, $\operatorname{Bour}(\sigma)=12$ and $c_{3}({\mathcal{T}}_{\sigma})=32$. Moreover, ${\mathcal{T}}_{\sigma}$ admits a minimal free resolution of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
given by the Buchsbaum–Rim complex.
2.2 Plane curves as pairs of surfaces
Let $g\in\kappa[x_{0},x_{1},x_{2}]$ be a square-free polynomial, so that the projective curve $X=V(g)\subset\mathbb{P}^{2}$ is reduced with isolated singularities, and the algebraically independent sequence $\sigma=(x_{3},g)$, denoting the plane $H=V(x_{3})\simeq\mathbb{P}^{2}$. We consider $S=V(g)\subset\mathbb{P}^{3}$ as a projective surface whose singular locus consists of possibly non-reduced lines, with the same multiplicity at each corresponding intersection point with $H$. Denote $Z_{g}\doteq\operatorname{Sing}(S)$ the associated singular scheme, given as the zero locus of the Jacobian ideal of $g$.